OpenStudy (anonymous):

Simplify the fraction 8/16 show your work

OpenStudy (anonymous):

8 -- 16

OpenStudy (jackellyn):

You find the biggest number the numerator and denominator have in common and divide by it. For example. 10/20 10 is the biggest number they have in common 10/10=1 ------ 20/10=2 So the answer for my example is 1/2. What is the biggest number 8 and 16 have in common?

OpenStudy (anonymous):

8 = 2 x 2 x 2 16 = 2 x 2 x 2 x 2 how many 2s do they have in common?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

2 x 2 x 2 = 8 then isn't 8 the biggest number both 8 and 16 have in common?

OpenStudy (jackellyn):

Exactly, like @nphuongsun93 says 8 is the biggest number that goes into both the numerator and the denominator so if you divide the top and bottom by 8 it will give you the answer.

OpenStudy (anonymous):

so my answers .50? thats what someone else said :o

OpenStudy (jackellyn):

8/8= 1 ---- -- 16/8= 2 so it is 1/2 which is 0.50.

OpenStudy (anonymous):

Thank you guys ssssoooo much!!!!

OpenStudy (jackellyn):

No problem @WhiteFlower , thanks for your help @nphuongsun93 :)